# Online Estimation with Rolling Validation: Adaptive Nonparametric Estimation with Streaming Data

### v3

@@ G.1 Regarding Batch Sieve Estimators
$E\left[\left\{\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\right\}^{2}\mid F^{j}\right]$ $=E\left[\left\{\hat{f}_{i-1}\left(X_{i};Z_{i}\right)-\hat{f}_{i-1}^{\prime}\left(X_{i};Z_{i}^{j}\right)\right\}^{2}\mid F^{j}\right]$ (145)
$=E\left[\left\{\left\langle\hat{\beta}_{i-1}^{\wedge},\bm{\phi}_{i}\right\rangle-\left\langle\hat{\beta}_{i-1}^{\prime\wedge},\bm{\phi}_{i}\right\rangle\right\}^{2}\mid F^{j}\right]$
$=E\left[\left(\hat{\eta}_{i-1}^{\wedge\top}\bm{\phi}_{i}\right)^{2}\mid F^{j}\right]$
$=E\left[\left(\hat{\theta}_{i-1}^{\wedge\top}W_{i}\right)^{2}\mid F^{j}\right]$
$=E\left[\hat{\theta}_{i-1}^{\top}E\left[W_{i}W_{i}^{\top}\right]\hat{\theta}_{i-1}\mid F^{j}\right]$
$\lesssim\lambda_{\max}\left(E\left[W_{i}W_{i}^{\top}\right]\right)i^{4\omega\zeta-2}.$
To finalize the proof, we just need to bound the minimal (maximal) eigenvalue of matrix $E[\bm{\phi}_{i}\bm{\phi}_{i}^{\top}]$ ($E[W_{i}W_{i}^{\top}]$). We denote the density of $X$ (with respect to the measure $\nu$) as $p_{X}$. By our assumption there exists some positive constants $\bar{p}_{X},\underline{p}_{X}$ such that $0<\underline{p}_{X}\leq p_{X}(x)\leq\bar{p}_{X}<\infty$. Let $v$ denote any eigenvector of $E[\bm{\phi}_{i}\bm{\phi}_{i}^{\top}]$, we have:
@@
So we know $\lambda_{min}(E[\bm{\phi}_{i}\bm{\phi}_{i}^{\top}])\geq\underline{p}_{X}$. Similarly, let $u$ denote any eigenvector of $E[W_{i}W_{i}^{\top}]$:
$u^{\top}E[W_{i}W_{i}^{\top}]u=E\left[\left(\sum_{k=1}^{J_{i}}k^{-\omega}\phi_{k}(X_{i})u_{k}\right)^{2}\right]\lesssim\bar{p}_{X}\|u\|^{2}.$ (147)
So we know $\lambda_{max}(E[W_{i}W_{i}^{\top}])\lesssim\bar{p}_{X}$. Combine these eigenvalue results with (145), we conclude that
$E\left[\left\{\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\right\}^{2}\mid F^{j}\right]\lesssim i^{4\omega\zeta-2}$
@@ G.4 A Technical Lemma
Our readers can take $a=0$ for the first pass which alleviates the technical complexity. For the ease of notation, we use $E_{j}$ to denote the conditional expectation $E[\cdot|F^{j}]$.
@@
$E_{k-1}\left[\|\eta_{k}\|^{2}/\|\eta_{i}\|^{2}\right]$ $=E_{k-1}\left[\frac{\|\eta_{k-1}\|^{2}}{\|\eta_{i}\|^{2}}\frac{\|\eta_{k}\|^{2}}{\|\eta_{k-1}\|^{2}}\right]=\frac{\|\eta_{k-1}\|^{2}}{\|\eta_{i}\|^{2}}E_{k-1}\left[\frac{\|\eta_{k}\|^{2}}{\|\eta_{k-1}\|^{2}}\right]$ (150)
@@
$E_{i}[\|\eta_{k}\|/\|\eta_{i}\|]$ $\leq\sqrt{E_{i}[\|\eta_{k}\|^{2}/\|\eta_{i}\|^{2}]}\leq\exp\left(-A\sum_{l=i}^{k-1}l^{-a}/2\right)$ (152)
$\Rightarrow E_{i}[\|\eta_{k}\|]$ $\leq\exp\left(-A\sum_{l=i}^{k-1}l^{-a}/2\right)\|\eta_{i}\|.$
@@
$n^{2}E_{j}[\|\hat{\eta}_{n}\|^{2}]$ $=E_{j}\left[\left\|\sum_{i=j}^{n}\eta_{i}\right\|^{2}\right]\quad\text{ ($\eta_{i}=0$ for $i<j$)}$ (153)
$\leq 2\sum_{i=j}^{n}\sum_{k=i}^{n}E_{j}[\eta_{k}^{\top}\eta_{i}]$
$\leq 2\sum_{i=j}^{n}\sum_{k=i}^{n}E_{j}[\|\eta_{k}\|\|\eta_{i}\|]$
$\leq 2\sum_{i=j}^{n}\sum_{k=i}^{n}\exp\left(-A\sum_{l=i}^{k-1}l^{-a}/2\right)E_{j}[\|\eta_{i}\|^{2}]$
$\stackrel{{\scriptstyle(I)}}{{\leq}}C(a,A)\sum_{i=j}^{n}i^{a}E_{j}[\|\eta_{i}\|^{2}]$
@@
Step $(I)$ and $(II)$ are technical, we will present them below.
@@ H.2 Numerical Comparison using Larger Exponents $\xi$
Figure H.1: Model selection results of Example 7.1. (A) the true MSE of the candidate models; (B)-(D), average rankings of the models at different sample sizes over 500 repetitions, according to wRV, with weight exponent $\xi=1,3,4$, respectively.

### v4

@@ G.1 Regarding Batch Sieve Estimators
$\mathbb{E}\left[\left\{\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\right\}^{2}\mid F^{j}\right]$ $=\mathbb{E}\left[\left\{\hat{f}_{i-1}\left(X_{i};Z_{i}\right)-\hat{f}_{i-1}^{\prime}\left(X_{i};Z_{i}^{j}\right)\right\}^{2}\mid F^{j}\right]$ (52)
$=\mathbb{E}\left[\left\{\left\langle\hat{\beta}_{i-1}^{\wedge},\bm{\phi}_{i}\right\rangle-\left\langle\hat{\beta}_{i-1}^{\prime\wedge},\bm{\phi}_{i}\right\rangle\right\}^{2}\mid F^{j}\right]$
$=\mathbb{E}\left[\left(\hat{\eta}_{i-1}^{\wedge\top}\bm{\phi}_{i}\right)^{2}\mid F^{j}\right]$
$=\mathbb{E}\left[\left(\hat{\theta}_{i-1}^{\wedge\top}W_{i}\right)^{2}\mid F^{j}\right]$
$=\mathbb{E}\left[\hat{\theta}_{i-1}^{\top}\mathbb{E}\left[W_{i}W_{i}^{\top}\right]\hat{\theta}_{i-1}\mid F^{j}\right]$
$\lesssim\lambda_{\max}\left(\mathbb{E}\left[W_{i}W_{i}^{\top}\right]\right)i^{4\omega\zeta-2}.$
To finalize the proof, we just need to bound the minimal (maximal) eigenvalue of matrix $\mathbb{E}[\bm{\phi}_{i}\bm{\phi}_{i}^{\top}]$ ($\mathbb{E}[W_{i}W_{i}^{\top}]$). We denote the density of $X$ (with respect to the measure $\nu$) as $p_{X}$. By our assumption there exists some positive constants $\bar{p}_{X},\underline{p}_{X}$ such that $0<\underline{p}_{X}\leq p_{X}(x)\leq\bar{p}_{X}<\infty$. Let $v$ denote any eigenvector of $\mathbb{E}[\bm{\phi}_{i}\bm{\phi}_{i}^{\top}]$, we have:
@@
So we know $\lambda_{min}(\mathbb{E}[\bm{\phi}_{i}\bm{\phi}_{i}^{\top}])\geq\underline{p}_{X}$. Similarly, let $u$ denote any eigenvector of $\mathbb{E}[W_{i}W_{i}^{\top}]$:
$u^{\top}\mathbb{E}[W_{i}W_{i}^{\top}]u=\mathbb{E}\left[\left(\sum_{k=1}^{J_{i}}k^{-\omega}\phi_{k}(X_{i})u_{k}\right)^{2}\right]\lesssim\bar{p}_{X}\|u\|^{2}.$
So we know $\lambda_{max}(\mathbb{E}[W_{i}W_{i}^{\top}])\lesssim\bar{p}_{X}$. Combining these eigenvalue results with (52), we conclude that
$\mathbb{E}\left[\left\{\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\right\}^{2}\mid F^{j}\right]\lesssim i^{4\omega\zeta-2}$
@@ G.4 A Technical Lemma
Our readers can take $a=0$ for the first pass which alleviates the technical complexity. For the ease of notation, we use $\mathbb{E}_{j}$ to denote the conditional expectation $\mathbb{E}[\cdot|F^{j}]$.
@@
$\mathbb{E}_{k-1}\left[\|\eta_{k}\|^{2}/\|\eta_{i}\|^{2}\right]$ $=\mathbb{E}_{k-1}\left[\frac{\|\eta_{k-1}\|^{2}}{\|\eta_{i}\|^{2}}\frac{\|\eta_{k}\|^{2}}{\|\eta_{k-1}\|^{2}}\right]=\frac{\|\eta_{k-1}\|^{2}}{\|\eta_{i}\|^{2}}\mathbb{E}_{k-1}\left[\frac{\|\eta_{k}\|^{2}}{\|\eta_{k-1}\|^{2}}\right]$
@@
$\mathbb{E}_{i}[\|\eta_{k}\|/\|\eta_{i}\|]$ $\leq\sqrt{\mathbb{E}_{i}[\|\eta_{k}\|^{2}/\|\eta_{i}\|^{2}]}\leq\exp\left(-A\sum_{l=i}^{k-1}l^{-a}/2\right)$
$\Rightarrow\mathbb{E}_{i}[\|\eta_{k}\|]$ $\leq\exp\left(-A\sum_{l=i}^{k-1}l^{-a}/2\right)\|\eta_{i}\|.$
@@
$n^{2}\mathbb{E}_{j}[\|\hat{\eta}_{n}\|^{2}]$ $=\mathbb{E}_{j}\left[\left\|\sum_{i=j}^{n}\eta_{i}\right\|^{2}\right]\quad\text{ ($\eta_{i}=0$ for $i<j$)}$ (54)
$\leq 2\sum_{i=j}^{n}\sum_{k=i}^{n}\mathbb{E}_{j}[\eta_{k}^{\top}\eta_{i}]$
$\leq 2\sum_{i=j}^{n}\sum_{k=i}^{n}\mathbb{E}_{j}[\|\eta_{k}\|\|\eta_{i}\|]$
$\leq 2\sum_{i=j}^{n}\sum_{k=i}^{n}\exp\left(-A\sum_{l=i}^{k-1}l^{-a}/2\right)\mathbb{E}_{j}[\|\eta_{i}\|^{2}]$
$\stackrel{{\scriptstyle(I)}}{{\leq}}C(a,A)\sum_{i=j}^{n}i^{a}\mathbb{E}_{j}[\|\eta_{i}\|^{2}]$
@@
Steps $(I)$ and $(II)$ are technical; we will present them below.
@@ H.2 Numerical Comparison using Larger Exponents $\xi$
Figure H.1: Model selection results of Example 7.1. (A) the true MSE of the candidate models; (B)-(D), average rankings of the models at different sample sizes over 500 repetitions, according to wRV, with weighting exponent $\xi=1,3,4$, respectively.